n find I (using mesh analysis) Į - lojn 2016 v ojn 2 long Iy=1.5 I
Answer #1

Given circuit is -lojsh In=1.5I 6jr 2010 12x6j 12+6; > 72j ciz+6;) Ix=15I 2040° using mesh anaysla i, x(-10s ) - z (, -iz) -2010+ loj I+ 130I; 12+6) =(10j+ 130; - -2060° 12 +6j f22.3L74.14) = -20L6° 22.8474.74 0.877 05.26